# Adapter-Enhanced Semantic Prompting for Continual Learning

### v2

@@ III-C Multimodal Prompt
For the $t$-th task dataset, $class1$, $class2,\dots,classN_{inc}$ are substituted with the actual class names from $Y(t)$. We then employ the text encoder of the pre-trained large language model BERT [26] as a semantic feature extractor to transform $R_{lan}$ into a text embedding:
@@
III-D Integrated Query-Key Matching
@@
To optimize the alignment of the query feature with its corresponding key while repelling other keys, we employ a cross-entropy loss function [54] as the multi-key loss:
@@
where $\ell$ represents the softmax-transformed version of $\hat{\ell}$, with $\ell_{i}$ corresponding to its i-th value. More details are given in Section IV-D.
@@
where $T\in\left\{1:\mathcal{T}\right\}$ represents the number of tasks already trained. Then, we adopt a voting strategy [55] to select the final task prompts. If all outputs differ, the final prompts are chosen by $P_{1}$.
III-E Semantic Adapter
In Section III-C, we advocate for the utilization of multimodal prompts to augment the model’s input processing capabilities. However, the representation spaces of semantic and visual information are substantially dissimilar, and the exclusive reliance on the frozen attention modules within the ViT architecture may prove inadequate for the effective integration of multimodal data, thereby potentially limiting the model’s generalization prowess. To alleviate this challenge, we propose the integration of semantic adapters into pre-trained models, facilitating robust visual-semantic interaction and the acquisition of more transferable feature representations.
@@ III-E Semantic Adapter
Building upon the innovative query-key matching mechanism introduced in Section III-D, which markedly improves the accuracy of task matching, we employ task-specific adapter parameters to mitigate the forgetting issue that arises from parameter updates. Upon the introduction of a new task, the adapter learns new parameters while preserving the existing parameters for each preceding task. During the inference phase, the query-key matching process is instrumental in selecting both the most pertinent prompts and the corresponding semantic adapters, thereby facilitating efficient and lightweight feature adaptation.
III-F Loss Functions
Semantic Contrast Loss. To ensure the semantic consistency and bolster the stability of our models, we introduce a novel Semantic Contrast Loss. For each $i$-th label in ${Y}(t)$ the corresponding class names are systematically structured into class-level semantic expressions, denoted as $R_{sem,i}$. These expressions are subsequently encoded into a semantic space using the BERT model [26], facilitating a robust representation:
@@ III-F Loss Functions
IV Experiments
IV-A Experimental Settings
@@
Datasets and Protocols. To facilitate a comprehensive comparison of various continual learning methodologies, we perform a series of experiments across three benchmark datasets, each characterized by distinct challenges, within the context of class-incremental learning [52, 47]. In this setting, task identities are obscured during the testing phase, thereby simulating a more realistic scenario. The datasets are delineated as follows:
@@
  • ImageNetR [56] is a dataset comprising 30,000 images across 200 classes derived from ImageNet. Each class features images in diverse styles, including art, cartoons, graffiti, and particularly challenging examples from the original ImageNet. This variety introduces a significant domain shift, making it especially difficult for models trained on standard datasets.
@@
  • CIFAR-100 [57] consists of 60,000 color images at a resolution of 32×32 pixels, categorized into 100 classes, with 500 training images and 100 test images per class. It is a widely recognized benchmark in the continual learning community.
@@
  • ImageNetA [58] is a real-world dataset containing 7,500 unmodified, naturally occurring images from 200 ImageNet classes. These images were specifically selected because they are misclassified by ResNet models, posing a significant challenge for machine learning models. Additionally, the dataset exhibits substantial class imbalance, with certain categories containing only a small number of samples.
TABLE I: Comparison of performance on Split-ImageNetR, CIFAR100, and Split-ImageNetA under the 10-task setting. The first four rows display the prompt-based methods, the following two rows present the adapter-based methods, and the proposed AESP is shown in the last row. The best result is marked in bold, and the second-best result is underlined.
Method Split-ImageNetR CIFAR100 Split-ImageNetA
Last-acc↑ Avg-acc↑ FF↓ Last-acc↑ Avg-acc↑ FF↓ Last-acc↑ Avg-acc↑ FF↓
L2P[20] 69.11±0.42 75.61±0.80 7.93±0.05 82.44±0.56 88.00±0.97 7.04±1.48 44.29±0.74 53.93±1.05 9.98 ±0.15
DualPrompt[21] 71.50±0.23 76.62±0.68 6.03±0.88 83.07±0.55 88.41±1.19 6.22±1.43 46.65±0.19 58.06±0.91 12.69±1.11
CODA-P[46] 75.17±0.23 80.59±0.68 7.07±0.86 86.19±0.36 90.97±1.19 6.70±0.88 51.90±0.71 62.23±1.42 9.69±0.67
Cprompt[47] 76.93±0.55 82.45±0.80 5.44 ±0.41 87.62±0.16 92.33±0.28 5.20±0.32 55.06±1.19 66.25±1.95 12.14±0.78
EASE[51] 76.16±0.25 82.85±0.46 7.45±0.34 88.12±0.56 92.49±0.75 5.83±0.60 54.16±0.20 65.98±0.78 13.05±0.82
SSIAT[50] 79.90 ±0.43 84.06 ±0.64 5.91±0.43 91.43 ±0.16 94.27 ±0.74 4.63 ±0.48 61.29 ±1.10 69.91 ±1.54 15.67±0.68
AESP(ours) 82.08±0.62 86.08±0.52 4.48±0.25 92.27±0.21 95.11±0.01 2.95±0.27 64.54±0.36 72.16±0.43 9.22±0.54
@@
TABLE II: Experimental results under 20-task setting on Split-ImageNetR and Split-ImageNetA dataset. The best result is marked in bold, and the second-best result is underlined.
@@
L2P[20] 65.88±0.33 73.18±1.37 37.16±0.36 47.95±0.43
DualPrompt[21] 67.80±0.51 73.82±1.45 39.76±0.24 53.67±0.72
CODA-P[46] 72.41±0.37 77.98±1.01 43.78±0.57 56.03±0.69
Cprompt[47] 74.31±0.79 81.18±0.15 52.29±0.81 64.02±1.26
EASE[51] 74.09±0.36 81.51±0.47 42.42±1.29 57.36±1.26
SSIAT[50] 78.32 ±0.36 82.99 ±0.68 58.45 ±2.23 68.03 ±2.18
AESP(ours) 79.83±0.55 84.5±0.56 61.39±0.49 70.66±0.86
Implementation Details. Following Cprompt [47], we utilize the Stochastic Gradient Descent (SGD) optimizer [59] with a momentum of 0.9 and an initial learning rate of 0.01. Our experiments were performed on an NVIDIA RTX 3090 GPU, using a batch size of 24 images per iteration. The learning rate is reduced to zero following a cosine annealing schedule. We trained our model on ImageNetR for 14 epochs and on CIFAR100 and ImageNetA for 20 epochs.
Evaluation Metrics. We assessed our model using three widely adopted metrics in continual learning [20, 21, 46, 6, 47]. Specifically, we measure the average prediction accuracy across all classes after the final training session, referred to as Last-acc, and the average accuracy across all sessions, denoted as Avg-acc. We also compute the forgetting score FF, following the approach outlined in [21, 46], to evaluate the performance degradation of previous tasks over time. To ensure the reliability and statistical significance of our findings, we conducted each experiment three times, employing different random seeds to vary the dataset order. We report the mean performance metrics of these runs, accompanied by their standard deviations.
IV-B Comparison with state-of-the-arts
@@
To thoroughly assess the efficacy of our proposed method, we conduct a comparative analysis against leading prompt-based and adapter-based state-of-the-art (SOTA) continual learning techniques, including L2P [20], DualPrompt [21], CODA-P [46], CPrompt [47], EASE [51], and SSIAT [50]. These methodologies were chosen due to their representative status at the cutting edge of their respective domains, providing a robust benchmark for evaluation. To ensure a fair comparison, we replicate the optimal results of these methods under identical experimental conditions. The results presented are the average of three independent runs, each with a unique random seed, to ensure statistical reliability. As demonstrated in Table I, the proposed AESP method consistently outperformed state-of-the-art (SOTA) methods across all datasets. Our method shows a substantial improvement over prompt-based approaches. Notably, while Cprompt has achieved significant progress beyond earlier prompt-based techniques, AESP surpasses it on ImageNetR, with a 5.15% higher Last-acc, a 3.63% higher Avg-acc, and a 0.96% lower forgetting rate. Similarly, on CIFAR100, AESP outperforms Cprompt by 4.65% in Last-acc and 2.78% in Avg-acc. Moreover, AESP achieves approximately a 2.25% lower forgetting rate.
@@
Furthermore, although the latest adapter-based methods have outperformed prompt-based approaches, our AESP method, which combines both adapters and prompts, exceeds these adapter-based techniques. Specifically, on ImageNetR, AESP achieves gains of 2.18% in Last-acc and 2.02% in Avg-acc over the leading adapter-based method SSIAT. While our advantage in Last-acc over prompt-based methods is notable, our advantage in Avg-acc over adapter-based methods is even more pronounced. This stems from the stronger resistance to the catastrophic forgetting of adapter-based methods. By combining adapters and prompts, AESP effectively reduces the forgetting rate by about 1.43% on ImageNetR and 1.68% on CIFAR100 compared to SSIAT.
@@
ImageNetA poses a unique challenge, consisting of samples misclassified by ResNet models due to their adversarial or anomalous nature. However, our method uses semantic information to improve the generalization of image features, improving the classification accuracy of these hard samples. AESP achieves a Last-acc of 64.54% and an Avg-acc of 72.16%, reflecting gains of about 3.25% and 2.25%, respectively, over SSIAT. A detailed comparison of different continual learning methods for each task is provided in Figure 3. It shows that our approach achieves excellent performance on each task and has a lower forgetting rate.
@@
Additionally, we assessed AESP under a 20-task setting to evaluate its performance across scenarios involving long sequences of tasks, as shown in Table II. Under the 20-task setting, AESP achieves a Last-acc of 79.83% on ImageNetR and 61.36% on ImageNetA, sustaining leading performance. These Last-acc results exceed SSIAT by 1.51% and Cprompt by up to 5.8%. A similar trend is observed for Avg-acc, indicating that our method maintains nice stability and strong resistance to forgetting when dealing with longer task sequences.
@@
TABLE III: Ablation Study on Split-ImageNetR and CIFAR100 dataset. We removed the Semantic Adapter (S-Adapter), Semantic Prompt (S-Prompt), and Integrated query-key matching (IQKM) from the model separately to verify the validity of each component.
@@
IV-C Ablation Study
We conducted ablation experiments under the 10-task setting on two datasets to evaluate the effectiveness of three components: semantic adapters, semantic prompts, and integrated query-key matching (IQKM). The results are shown in Table III. The results show that when the semantic adapters are removed from the ViT layers, the Last-acc decreased by 1.63% and 0.87% on ImageNetR and CIFAR100, respectively, while the forgetting score FF remained at a similar level. This indicates that the adapters positively contribute to the model’s recognition accuracy without significantly affecting forgetting.
@@ IV-C Ablation Study
Furthermore, when the IQKM module was replaced with the Multi-Key matching mechanism for query-key matching, significant performance drops occurred across both datasets. Specifically, Last-acc decreased by 7.69% and 7.73% on ImageNetR and CIFAR100, respectively, and the forgetting score FF increased by 3.77% and 3.86%. This substantial decline indicates the importance of IQKM in achieving effective task selection, as detailed in Section IV-D. Inaccurate task selection leads to the use of inappropriate prompts and adapters, resulting in reduced accuracy and increased catastrophic forgetting.
IV-D Detailed Analysis
@@
Effectiveness of IQKM. To enhance prompt selection accuracy, we propose the Integrated Query-Key Matching (IQKM), as detailed in Section III-D. This approach amalgamates Multi-Key matching, Entropy-based matching, and Prototype matching through a consensus voting mechanism to identify the appropriate task. To ascertain the efficacy of IQKM, we conducted an evaluation of task selection accuracy across various methods using the ImageNetR dataset. As illustrated in Figure 4, we utilize four heatmaps to assess the individual contributions of each strategy. Each heatmap delineates the relationship between the query task (ground-truth task) and the task selected by the model. It is shown that all three matching approaches have high task selection errors. In contrast, IQKM demonstrates a marked improvement in task selection accuracy. Furthermore,we quantitatively measured the task selection accuracy across all three datasets, with the results presented in Table IV. It shows that IQKM achieves 98.55%, 94.48%, and 93.55% on Split-ImageNetR, CIFAR100, and Split-ImageNetA, respectively, significantly outperforming other approaches, which benefits the image recognition.
TABLE IV: The task selection accuracy of each prompt-key matching on three datasets.
Method ImageNetR CIFAR100 ImageNetA
Multi-Key Matching 64.77 58.55 52.47
Entropy-based Matching 42.77 47.52 41.34
Prototype Matching 66.98 75.17 56.95
IQKM 98.55 94.48 93.55
Detailed Analysis of Semantic Prompt. To improve the generalization of image features, we utilize semantic prompts encoded by a text encoder to enrich the input information, as described in III-C. To validate the effectiveness of semantic embedding, we compare our AESP method with a conventional prompt-based approach, wherein the learnable prompt is randomly initialized and subsequently updated during training. The results, presented in Table V for the ImageNet-R dataset, indicate that our method achieves a forget rate of 4.48%, which is significantly lower than that of using trainable prompts. Furthermore, our approach yields a Last-acc of 82.08% and an Avg-acc of 86.08%, underscoring the semantic prompt’s efficacy in alleviating catastrophic forgetting and preserving the pre-trained network’s performance. In addition, we introduce a semantic contrastive loss to ensure semantic consistency. To assess its impact, we conducted an experiment replacing the proposed semantic contrast loss with the commonly used cross-entropy loss, with the results displayed in the second line of Table V. The performance decline observed in the absence of semantic contrastive loss substantiates its critical role in maintaining model performance.
TABLE V: Detail analysis of Semantic Prompt on 10-task continual learning of ImagenetR.
@@
Influence of text encoder. In our prior experiments, we employed the BERT text encoder to transform semantic expressions into a semantic space. To investigate the impact of different text encoders on our AESP method, we further evaluated its performance using two additional pre-trained text transformer models: CLIP [25] and Sentence-BERT [60]. CLIP is designed to learn aligned embeddings for images and text in a shared space, whereas Sentence-BERT creates sentence embeddings optimized for semantic similarity tasks through fine-tuning of BERT. The results for ImageNetR, presented in Table VI, reveal that our method exhibits resilience to the choice of text encoder. Notably, CLIP’s text encoder, trained on an extensive corpus of image-text pairs for visual-semantic alignment, achieves the highest Last-Acc score. Sentence-BERT, with its simpler semantic representation, yields a more modest performance improvement. Nonetheless, the consistent performance gains across different text encoders underscore the robustness of AESP to variations in text encoding mechanisms.
TABLE VI: Influence of different text encoders. We test our proposed method with CLIP, BERT, and Sentence-BERT text encoders on 10-task continual learning of ImagenetR and cifar100.
@@
CLIP [25] 82.27 4.67 92.43 3.22
Sen-BERT [60] 82.08 4.48 92.27 2.95
Bert [26] 81.81 4.47 92.4 2.97
Impact of semantic adapter location. The semantic adapter plays a crucial role in fusing visual and semantic information, thereby enhancing the generalization capabilities of image features. This adapter can be integrated into various layers of the Vision Transformer (ViT). To determine the optimal number and placement of semantic adapters, we conducted a series of experiments on the ImageNetR dataset within a 10-task setting. The results, as presented in Table VII, indicate that incorporating adapters into all layers of the ViT yields the highest performance. Furthermore, our findings reveal that placing adapters in shallow layers results in superior outcomes compared to deeper layers. This observation diverges from earlier research [9, 23], which primarily focused on adapter-based methods dealing with pure image information. In our approach, the integration of semantic information suggests that shallow layers may be more conducive to effective feature adaptation, highlighting a distinct advantage of our method in leveraging semantic context.
TABLE VII: The influence of Adapters for different layers. Experiments were performed on the Split-imagenetR dataset.

### v3

@@ 3.3 Multimodal Prompt
3.4 Integrated Query-Key Matching
@@
To optimize the alignment of the query feature with its corresponding key while repelling other keys, we employ a cross-entropy loss function Zhang and Sabuncu (2018) as the multi-key loss:
@@
where $\ell$ represents the softmax-transformed version of $\hat{\ell}$, with $\ell_{i}$ corresponding to its i-th value. More details are given in Section 4.4.
@@
where $T\in\left\{1:\mathcal{T}\right\}$ represents the number of tasks already trained. Then, we adopt a voting strategy Breiman (2001) to select the final task prompts. If all outputs differ, the final prompts are chosen by $P_{1}$.
3.5 Semantic Adapter
In Section 3.3, we advocate for the utilization of multimodal prompts to augment the model’s input processing capabilities. However, the representation spaces of semantic and visual information are substantially dissimilar, and the exclusive reliance on the frozen attention modules within the ViT architecture may prove inadequate for the effective integration of multimodal data, thereby potentially limiting the model’s generalization prowess. To alleviate this challenge, we propose the integration of semantic adapters into pre-trained models, facilitating robust visual-semantic interaction and the acquisition of more transferable feature representations.
@@ 3.5 Semantic Adapter
Building upon the innovative query-key matching mechanism introduced in Section 3.4, which markedly improves the accuracy of task matching, we employ task-specific adapter parameters to mitigate the forgetting issue that arises from parameter updates. Upon the introduction of a new task, the adapter learns new parameters while preserving the existing parameters for each preceding task. During the inference phase, the query-key matching process is instrumental in selecting both the most pertinent prompts and the corresponding semantic adapters, thereby facilitating efficient and lightweight feature adaptation.
3.6 Loss Functions
Semantic Contrast Loss.   To ensure the semantic consistency and bolster the stability of our models, we introduce a novel Semantic Contrast Loss. For each $i$-th label in ${Y}(t)$ the corresponding class names are systematically structured into class-level semantic expressions, denoted as $R_{sem,i}$. These expressions are subsequently encoded into a semantic space using the BERT model Devlin et al. (2019), facilitating a robust representation:
@@ 3.6 Loss Functions
[…]
Method Reference Split-ImageNetR
Last-acc↑ Avg-acc↑ FF↓
L2P CVPR 2022 Wang et al. (2022b) 69.11±0.42 75.61±0.80 7.93±0.05
DualPrompt ECCV 2022 Wang et al. (2022a) 71.50±0.23 76.62±0.68 6.03±0.88
CODA-Prompt CVPR 2023 Smith et al. (2023) 75.17±0.23 80.59±0.68 7.07±0.86
Cprompt CVPR 2024 Gao et al. (2024b) 76.93±0.55 82.45±0.80 5.44±0.41
EASE CVPR 2024 Zhou et al. (2024a) 76.16±0.25 82.85±0.46 7.45±0.34
SSIAT CVPR 2024 Tan et al. (2024) 79.38±0.59 83.63±0.43 -
ADAM IJCV 2025 Zhou et al. (2025) 72.87 79.39 -
Lora-DRS CVPR 2025 Liu and Chang (2025) 74.74±0.78 81.16±0.59 -
DIA CVPR 2025 Li et al. (2025) 79.03 85.61 -
AESP This Work 82.08±0.62 86.08±0.52 4.48±0.25
Table 2: Comparison of performance on CIFAR100 and Split-ImageNetA under the 10-task setting.
Method CIFAR100 Split-ImageNetA
Last-acc↑ Avg-acc↑ FF↓ Last-acc↑ Avg-acc↑ FF↓
L2P 82.44±0.56 88.00±0.97 7.04±1.48 44.29±0.74 53.93±1.05 9.98 ±0.15
DualPrompt 83.07±0.55 88.41±1.19 6.22±1.43 46.65±0.19 58.06±0.91 12.69±1.11
CODA-Prompt 86.19±0.36 90.97±1.19 6.70±0.88 51.90±0.71 62.23±1.42 9.69±0.67
Cprompt 87.62±0.16 92.33±0.28 5.20±0.32 55.06±1.19 66.25±1.95 12.14±0.78
EASE 88.12±0.56 92.49±0.75 5.83±0.60 54.16±0.20 65.98±0.78 13.05±0.82
SSIAT 91.35 ±0.26 94.35 ±0.60 - 62.43 ±1.63 70.83 ±1.63 -
ADAM 85.75 90.94 - 49.51 59.89 -
Lora-DRS 89.14±0.23 92.55±0.25 - - - -
DIA 90.8 94.29 - 61.69 71.58 -
AESP(ours) 92.27±0.21 95.11±0.01 2.95±0.27 64.54±0.36 72.16±0.43 9.22±0.54
4 Experiments
4.1 Experimental Settings
@@
Datasets and Protocols.    To facilitate a comprehensive comparison of various continual learning methodologies, we perform a series of experiments across three benchmark datasets, each characterized by distinct challenges, within the context of class-incremental learning  (Mittal et al., 2021; Gao et al., 2024b). In this setting, task identities are obscured during the testing phase, thereby simulating a more realistic scenario. The datasets are delineated as follows:
@@
  • ImageNetR  (Hendrycks et al., 2021a) is a dataset comprising 30,000 images across 200 classes derived from ImageNet. Each class features images in diverse styles, including art, cartoons, graffiti, and particularly challenging examples from the original ImageNet. This variety introduces a significant domain shift, making it especially difficult for models trained on standard datasets.
@@
  • CIFAR-100  (Krizhevsky and Hinton, 2009) consists of 60,000 color images at a resolution of 32×32 pixels, categorized into 100 classes, with 500 training images and 100 test images per class. It is a widely recognized benchmark in the continual learning community.
@@
  • ImageNetA  (Hendrycks et al., 2021b) is a real-world dataset containing 7,500 unmodified, naturally occurring images from 200 ImageNet classes. These images were specifically selected because they are misclassified by ResNet models, posing a significant challenge for machine learning models. Additionally, the dataset exhibits substantial class imbalance, with certain categories containing only a small number of samples.
Implementation Details.   Following Cprompt  (Gao et al., 2024b), we utilize the Stochastic Gradient Descent (SGD) optimizer  (Keskar and Socher, 2017) with a momentum of 0.9 and an initial learning rate of 0.01. Our experiments were performed on an NVIDIA RTX 3090 GPU, using a batch size of 24 images per iteration. The learning rate is reduced to zero following a cosine annealing schedule. We trained our model on ImageNetR for 14 epochs and on CIFAR100 and ImageNetA for 20 epochs.
Evaluation Metrics.   We assessed our model using three widely adopted metrics in continual learning  (Wang et al., 2022b, a; Smith et al., 2023; Wang et al., 2024; Gao et al., 2024b). Specifically, we measure the average prediction accuracy across all classes after the final training session, referred to as Last-acc, and the average accuracy across all sessions, denoted as Avg-acc. We also compute the forgetting score FF, following the approach outlined in  (Wang et al., 2022a; Smith et al., 2023), to evaluate the performance degradation of previous tasks over time. To ensure the reliability and statistical significance of our findings, we conducted each experiment three times, employing different random seeds to vary the dataset order. We report the mean performance metrics of these runs, accompanied by their standard deviations.
@@
[…]
@@
L2P 65.88±0.33 73.18±1.37 37.16±0.36 47.95±0.43
Dual-P 67.80±0.51 73.82±1.45 39.76±0.24 53.67±0.72
CODA-P 72.41±0.37 77.98±1.01 43.78±0.57 56.03±0.69
Cprompt 74.31±0.79 81.18±0.15 52.29±0.81 64.02±1.26
EASE 74.09±0.36 81.51±0.47 42.42±1.29 57.36±1.26
SSIAT 75.67 ±0.14 82.30 ±0.36 59.16 ±1.03 68.45 ±1.92
ADAM 70.47 77.29 49.57 60.53
Lora-DRS 74.80 ±0.73 80.69 ±0.75 - -
DIA 76.32 83.51 - -
AESP 79.83±0.55 84.5±0.56 61.39±0.49 70.66±0.86
4.2 Comparison with state-of-the-arts
@@
To thoroughly assess the efficacy of our proposed method, we conduct a comparative analysis against leading prompt-based and adapter-based state-of-the-art (SOTA) continual learning techniques, including L2P  (Wang et al., 2022b), DualPrompt  (Wang et al., 2022a), CODA-P  (Smith et al., 2023), CPrompt  (Gao et al., 2024b), EASE  (Zhou et al., 2024a), and SSIAT  (Tan et al., 2024). These methodologies were chosen due to their representative status at the cutting edge of their respective domains, providing a robust benchmark for evaluation. To ensure a fair comparison, we replicate the optimal results of these methods under identical experimental conditions. The results presented are the average of three independent runs, each with a unique random seed, to ensure statistical reliability. As demonstrated in Table 1 and Table 2, the proposed AESP method consistently outperformed state-of-the-art (SOTA) methods across all datasets. Our method shows a substantial improvement over prompt-based approaches. Notably, while Cprompt has achieved significant progress beyond earlier prompt-based techniques, AESP surpasses it on ImageNetR, with a 5.15% higher Last-acc, a 3.63% higher Avg-acc, and a 0.96% lower forgetting rate. Similarly, on CIFAR100, AESP outperforms Cprompt by 4.65% in Last-acc and 2.78% in Avg-acc. Moreover, AESP achieves approximately a 2.25% lower forgetting rate.
@@
Furthermore, although the latest adapter-based methods have outperformed prompt-based approaches, our AESP method, which combines both adapters and prompts, exceeds these adapter-based techniques. Specifically, on ImageNetR, AESP achieves gains of 2.45% in Last-acc and 2.7% in Avg-acc over the leading adapter-based method SSIAT. While our advantage in Last-acc over prompt-based methods is notable, our advantage in Avg-acc over adapter-based methods is even more pronounced. This stems from the stronger resistance to the catastrophic forgetting of adapter-based methods. By combining adapters and prompts, AESP effectively reduces the forgetting rate by about 2.97% on ImageNetR and 2.88% on CIFAR100 compared to SSIAT.
@@
ImageNetA poses a unique challenge, consisting of samples misclassified by ResNet models due to their adversarial or anomalous nature. However, our method uses semantic information to improve the generalization of image features, improving the classification accuracy of these hard samples. AESP achieves a Last-acc of 64.54% and an Avg-acc of 72.16%, reflecting gains of about 2.11% and 1.33%, respectively, over SSIAT. A detailed comparison of different continual learning methods for each task is provided in Figure  3. It shows that our approach achieves excellent performance on each task and has a lower forgetting rate.
@@
Additionally, we assessed AESP under a 20-task setting to evaluate its performance across scenarios involving long sequences of tasks, as shown in Table  3. Under the 20-task setting, AESP achieves a Last-acc of 79.83% on ImageNetR and 61.36% on ImageNetA, sustaining leading performance. These Last-acc results exceed SSIAT by 4.16% and Cprompt by up to 5.52%. A similar trend is observed for Avg-acc, indicating that our method maintains nice stability and strong resistance to forgetting when dealing with longer task sequences.
@@
Table 4: Ablation Study on Split-ImageNetR and CIFAR100 dataset. We removed the Semantic Adapter (S-Adapter), Semantic Prompt (S-Prompt), and Integrated query-key matching (IQKM) from the model separately to verify the validity of each component.
@@
4.3 Ablation Study
We conducted ablation experiments under the 10-task setting on two datasets to evaluate the effectiveness of three components: semantic adapters, semantic prompts, and integrated query-key matching (IQKM). The results are shown in Table  4. The results show that when the semantic adapters are removed from the ViT layers, the Last-acc decreased by 1.63% and 0.87% on ImageNetR and CIFAR100, respectively, while the forgetting score FF remained at a similar level. This indicates that the adapters positively contribute to the model’s recognition accuracy without significantly affecting forgetting.
@@ 4.3 Ablation Study
Furthermore, when the IQKM module was replaced with the Multi-Key matching mechanism for query-key matching, significant performance drops occurred across both datasets. Specifically, Last-acc decreased by 7.69% and 7.73% on ImageNetR and CIFAR100, respectively, and the forgetting score FF increased by 3.77% and 3.86%. This substantial decline indicates the importance of IQKM in achieving effective task selection, as detailed in Section  4.4. Inaccurate task selection leads to the use of inappropriate prompts and adapters, resulting in reduced accuracy and increased catastrophic forgetting.
4.4 Detailed Analysis
@@
Effectiveness of IQKM.   To enhance prompt selection accuracy, we propose the Integrated Query-Key Matching (IQKM), as detailed in Section 3.4. This approach amalgamates Multi-Key matching, Entropy-based matching, and Prototype matching through a consensus voting mechanism to identify the appropriate task. To ascertain the efficacy of IQKM, we conducted an evaluation of task selection accuracy across various methods using the ImageNetR dataset. As shown in Figure 4, we visualize the contribution of each strategy through four heatmaps that reveal the correlation between ground-truth tasks (vertical axis) and model-predicted task selections (horizontal axis). The diagonal patterns in the heatmaps for three matching approaches indicate substantial task selection errors, particularly evident in their scattered distributions across non-diagonal regions. In contrast, IQKM demonstrates a marked improvement in task selection accuracy, which ensures consistency between knowledge acquisition during training and its application in inference. This alignment between learned representations and task-specific utilization guarantees robust feature encoding stability when handling cross-dataset inputs, preserving task-discriminative patterns that enhance recognition reliability.
Detailed Analysis of Semantic Prompt.   To improve the generalization of image features, we utilize semantic prompts encoded by a text encoder to enrich the input information, as described in  3.3. To validate the effectiveness of semantic embedding, we compare our AESP method with a conventional prompt-based approach, wherein the learnable prompt is randomly initialized and subsequently updated during training. The results, presented in Table  5 for the ImageNet-R dataset, indicate that our method achieves a forget rate of 4.48%, which is significantly lower than that of using trainable prompts. Furthermore, our approach yields a Last-acc of 82.08% and an Avg-acc of 86.08%, underscoring the semantic prompt’s efficacy in alleviating catastrophic forgetting and preserving the pre-trained network’s performance. In addition, we introduce a semantic contrastive loss to ensure semantic consistency. To assess its impact, we conducted an experiment replacing the proposed semantic contrast loss with the commonly used cross-entropy loss, with the results displayed in the second line of Table  5. The performance decline observed in the absence of semantic contrastive loss substantiates its critical role in maintaining model performance.
Table 5: Detail analysis of Semantic Prompt on 10-task continual learning of ImagenetR.
@@
Influence of text encoder.    In our prior experiments, we employed the BERT text encoder to transform semantic expressions into a semantic space. To investigate the impact of different text encoders on our AESP method, we further evaluated its performance using two additional pre-trained text transformer models: CLIP  (Radford et al., 2021) and Sentence-BERT  (Reimers and Gurevych, 2019). CLIP is designed to learn aligned embeddings for images and text in a shared space, whereas Sentence-BERT creates sentence embeddings optimized for semantic similarity tasks through fine-tuning of BERT. The results for ImageNetR, presented in Table  6, reveal that our method exhibits resilience to the choice of text encoder. Notably, CLIP’s text encoder, trained on an extensive corpus of image-text pairs for visual-semantic alignment, achieves the highest Last-Acc score. Sentence-BERT, with its simpler semantic representation, yields a more modest performance improvement. Nonetheless, the consistent performance gains across different text encoders underscore the robustness of AESP to variations in text encoding mechanisms.
Table 6: Influence of different text encoders. We test our proposed method with CLIP, BERT, and Sentence-BERT text encoders on 10-task continual learning of ImagenetR and cifar100.
@@
CLIP 82.27 4.67 92.43 3.22
Sen-BERT 82.08 4.48 92.27 2.95
Bert 81.81 4.47 92.4 2.97
Impact of semantic adapter location.    The semantic adapter plays a crucial role in fusing visual and semantic information, thereby enhancing the generalization capabilities of image features. This adapter can be integrated into various layers of the Vision Transformer (ViT). To determine the optimal number and placement of semantic adapters, we conducted a series of experiments on the ImageNetR dataset within a 10-task setting. The results, as presented in Table  7, indicate that incorporating adapters into all layers of the ViT yields the highest performance. Furthermore, our findings reveal that placing adapters in shallow layers results in superior outcomes compared to deeper layers. This observation diverges from earlier research  (Gao et al., 2024a; Park et al., 2024), which primarily focused on adapter-based methods dealing with pure image information. In our approach, the integration of semantic information suggests that shallow layers may be more conducive to effective feature adaptation, highlighting a distinct advantage of our method in leveraging semantic context.
Table 7: The influence of Adapters for different layers. Experiments were performed on the Split-imagenetR dataset.